# Robust Learning in Bayesian Parallel Branching Graph Neural Networks: The Narrow Width Limit

### v1

@@ 4.2 Student-Teacher experiment on robust branch learning
We demonstrate this robust learning phenomenon and provide a first evidence of the equipartition conjecture with the student-teacher experiment setup introduced in the previous section. We use the contextual stochastic block model (CSBM) (Deshpande et al. (2018)) as the shared node input features, where the adjacency matrix is given by a stochastic block model with two blocks, and the node feature is generated with latent vectors corresponding to the two blocks (See B.1 for details).
Figure 2: Student readout norms squared as a function of network width. Left: Student branches break the GP symmetry as it goes to the narrow width limit. Middle, Right: Branch 0 and branch 1 readout norm squared for a range of $\sigma$ regularization values. Different regularization strengths will all converge to the same teacher readout norm values.
As shown in Figure 1(b), an extremely narrow student network learns the teacher’s branch readout norms very robustly, despite having a large variance; on the other hand, a much wider network fails to learn the teachers’ norms and approaches the GP limit, while having a smaller variance.
@@
Figure 2 shows the student branch norms as a function of the network width. The symmetry breaking of branch norms from the GP limit to the narrow width limit accompanies the convergence to learning teacher’s norms at narrow width for different $\sigma$’s, supporting Eq. 21.
@@
Using the mean predictor and variance from the theory, we can determine the generalization of individual branches (Appendix A.2) as a function of network width $N$, as shown in Figure 3. At narrow width, we expect individual branch to learn the teacher’s branch output $y_{l}$ independently, causing the bias to increase with network width. This is observed for both branches, with a transition from the narrow-width regime to the GP regime. The regularization strength $\sigma^{2}$ controls the transition window, with larger $\sigma$’s leading to sharper transitions. This aligns with our analysis of the entropic and energetic contributions, where larger $\sigma$ amplifies the distinction between the two terms. In contrast, the variance decreases with network width for small $\sigma$’s, resulting in a trade-off between the contributions of bias and variance to overall generalization performance, as shown clearly with $\sigma=0.5$ in Figure 4.
Figure 3: Student network bias and variance for individual branch as a function of network width for student-teacher tasks.
Figure 4: Overall generalization performance as a function of network width for student-teacher tasks.
@@ 5 BPB-GNN on Cora
We also conduct experiments on the benchmark dataset Cora with binary node regression, and observe a similar narrow-to-wide width transition for the bias term. We analyze the generalization performance for both network width $N$ and the number of branches used $L$. As shown in Figure 5, the bias increases with network width, transitioning to the GP regime. Generally, more branches improve performance.
Figure 5: Cora generalization performance vs. network width $N$ and branch number $L$. The accuracy is computed by turning the mean predictor into a class label using its sign.
Convergence of Branch Importance at Narrow Width
An interesting aspect of the BPB-GNN network is that the branch norms seem to converge regardless of the hyperparameters $\sigma$ and $L$, reflecting the natural branch importance of the dataset and task.
Figure 6: Cora branch norms vs. network width $N$, $L=6$.
As shown in Figure 6, the BPB-GNN with $L=6$ branches robustly learns the readout norms at narrow width, verifying the equipartition conjecture. The last branch of the BPB-GNN network has a larger contribution, reflecting the presence of higher-order convolutions in the Cora dataset. From a kernel perspective, increasing branches better distinguish the nodes, as shown in Figure 7. This could explain the selective turn-off of intermediate branches and the increased contribution of the last branch.
Figure 7: Branch kernel $K_{l}$ visualizations the first 8 branches on Cora.
Figure 8: Branch Importance vs. L at narrow width.
Additionally, the first two branches are most robustly learned at narrow width, as shown in Figure 8, where the branch norms converge for the first two branches even for BPB-GNNs with different $L$. This suggests that the branch importance, as reflected by the norms learned at narrow width, indicates the contribution of the bare data and the first convolution layer.
@@ 6 Discussion
The findings presented in this paper reveal that BPB-GNNs exhibit unique characteristics in the narrow width limit. Unlike the infinite-width limit where neural networks behave as Gaussian Process (GP) with task-independent kernels, narrow-width BPB-GNNs undergo significant kernel renormalization. This renormalization leads to symmetry breaking among the branches, resulting in more robust and differentiated learning. Our experiments demonstrate that narrow-width BPB-GNNs can retain and, in some cases, improve generalization performance compared to their wider counterparts, particularly in bias-limited scenarios where the regularization effects dominate. Additionally, the observed independence of readout norms from architectural hyperparameters suggests that narrow-width BPB-GNNs can capture the intrinsic properties of the data more effectively. These insights suggest potential new strategies for optimizing neural network architectures in practical applications, challenging the traditional emphasis on increasing network width for better performance. Although our work is focused on GNN architectures, our findings can be transferred to the transformer family of architectures, as the transformer with frozen attention is analogous to a fully connected GNN (Vaswani et al. (2017); Veličković et al. (2017)), and a Bayesian theory with kernel renormalization can be similarly developed as in Tiberi et al. (2024).
@@ 7 Conclusion
In conclusion, this paper introduces and investigates the concept of narrow width limits in Bayesian Parallel Branching Graph Neural Networks. Contrary to the common belief that wider networks inherently generalize better, our results indicate that BPB-GNNs with significantly narrower widths can achieve better or competitive performance. This is attributed to effective symmetry breaking and kernel renormalization in the narrow-width limit, which lead to robust learning. Our theoretical analysis, supported by empirical evidence, establishes a new understanding of how network architecture influences learning outcomes. This work provides a novel perspective on neural network design and suggests further research into optimizing network structures for specific learning tasks, especially in scenarios where model simplicity and computational efficiency are crucial.
Acknowledgments and Disclosure of Funding
@@ Appendix A Details on Theory of BPB-GNN
A.1 Kernel renormalization
@@
$H(W)=\frac{1}{2\sigma^{2}}\sum_{l=0}^{L-1}\mathrm{Tr}W_{l}^{T}W_{l}+\frac{1}{2}Y^{T}(K(W)+TI)^{-1}Y+\frac{1}{2}\log\det(K(W)+TI),$ (26)
@@
$K(W)=\frac{1}{L}\sum_{l}\frac{\sigma^{2}}{N}(H_{l}(W_{l})H_{l}(W_{l})^{T})|_{P}$ (27)
@@
$Z$ $=\int\prod_{l=0}^{L}dh_{l}du_{l}dt\exp\Big{(}it^{T}Y-\sum_{l}\frac{N}{2}\log(1+h_{l})+\sum_{l}\frac{N}{2\sigma^{2}}u_{l}h_{l}-\frac{1}{2}t^{T}\Big{(}\sum_{l}\frac{1}{L}u_{l}K_{l}+TI\Big{)}t\Big{)}$
$=\int\prod_{l=0}^{L}dh_{l}du_{l}\exp\Big{(}-\sum_{l}\frac{N}{2}\log(1+h_{l})+\sum_{l}\frac{N}{2\sigma^{2}}u_{l}h_{l}\frac{1}{2}Y^{T}\Big{(}\sum_{l}\frac{1}{L}u_{l}K_{l}+TI\Big{)}^{-1}Y\Big{)}$ (28)
@@
$H_{eff}=S+E,$ (32)
where we call $S$ the entropic term
$S=-\sum_{l}\frac{N}{2}\log u_{l}+\sum_{l}\frac{N}{2\sigma^{2}}u_{l}$ (33)
and $E$ the energetic term
$E=\frac{1}{2}Y^{T}(\sum_{l}\frac{1}{L}u_{l}K_{l}+TI)^{-1}Y+\frac{1}{2}\log\det(\sum_{l}\frac{1}{L}u_{l}K_{l}+TI)$ (34)
@@
$N(1-\frac{u_{l}}{\sigma^{2}})=-Y^{T}(K+TI)^{-1}\frac{u_{l}K_{l}}{L}(K+TI)^{-1}Y+\mathrm{Tr}[K^{-1}\frac{u_{l}K_{l}}{L}],$ (37)
@@
$N(1-\frac{u_{l}}{\sigma^{2}})=-Y^{T}K^{-1}\frac{u_{l}K_{l}}{L}K^{-1}Y+\mathrm{Tr}[K^{-1}\frac{u_{l}K_{l}}{L}]$ (40)
A.2 Predictor statistics and generalization
@@
$Z(\eta_{1},\dots,\eta_{L})=\int D\Theta\exp{\left\{-\frac{\beta}{2}\sum_{\mu}(f^{\mu}(G;\Theta)-y^{\mu})^{2}+\sum_{l}i\eta_{l}\frac{1}{\sqrt{NL}}\sum_{i}a^{(l)}_{i}h_{i}^{(l),\nu}(G,W_{l})-\frac{T}{2\sigma^{2}}\Theta^{T}\Theta\right\}}$ (41)
@@
$\langle y_{l}^{\nu}(G)\rangle=\partial_{i\eta_{l}}\log Z\big{|}_{\vec{\eta}=0}$ (42)
$\langle\delta y_{l,\nu}^{2}(G)\rangle=\partial^{2}_{i\eta_{l}}\log Z\big{|}_{\vec{\eta}=0}$ (43)
@@
$Z(\eta_{1},\dots,\eta_{L})$ $=\int\Pi_{l}du_{l}\,\exp{\left\{\sum_{l}\left(\frac{N}{2}\log u_{l}-\frac{N}{2\sigma^{2}}u_{l}\right)\right.}$
@@
$\langle y_{l}^{\nu}\rangle=\frac{u_{l}k_{l,\nu}^{T}}{L}(K+TI)^{-1}Y$ (47)
@@
$\langle\delta y_{l,\nu}^{2}\rangle=\frac{u_{l}K_{l}^{\nu,\nu}}{L}-\frac{u_{l}k_{l,\nu}^{T}}{L}(K+TI)^{-1}\frac{u_{l}k_{l,\nu}}{L}$ (48)
The predictor statistics of the overall readout $y=\sum_{l}y_{l}$ is given by:
$\langle y^{\nu}(G)\rangle=\sum_{l}\frac{u_{l}k_{l,\nu}^{T}}{L}(K+TI)^{-1}Y=k_{\nu}^{T}(K+TI)^{-1}Y$ (49)
$\langle\delta y(G)_{\nu}^{2}\rangle=\sum_{l}u_{l}K_{l}^{\nu,\nu}-\sum_{l,l^{\prime}}u_{l}k_{l,\nu}^{T}(K+TI)^{-1}u_{l^{\prime}}k_{l^{\prime},\nu}=K_{\nu,\nu}-k_{\nu}^{T}(K+TI)^{-1}k_{\nu}$ (50)
A.3 Statistics of branch readout norms

### v2

@@ 4.2 Student-Teacher experiment on robust branch learning
We demonstrate this robust learning phenomenon and provide a first evidence of the equipartition conjecture with the student-teacher experiment setup introduced in the previous section. We use the contextual stochastic block model (CSBM) (Deshpande et al. (2018)) to generate the graph, where the adjacency matrix is given by a stochastic block model with two blocks, and the node feature is generated with latent vectors corresponding to the two blocks (Appendix B.1). Both student and teacher network has $L=2$ branches. We calculate the order parameters $u_{l}$’s for the student network with the saddle point equations, as well as perform Hamiltonian Monte Carlo (HMC) sampling (Appendix B.3) from the posterior distribution of the student’s network weights to generate a distribution of student readout norms squared $a_{l}^{2}$. We compare the values of the student branch readout norms to their corresponding teacher branches for the student network with either narrow or wide width. As shown in Figure 1(b), an extremely narrow student network ($N=4$) learns the teacher’s branch readout norms very robustly, while a much wider network ($N=1024$) fails to learn the teachers’ norms and approaches the GP ($N\to\infty$) limit where the two branches are indistinguishable.
Figure 2: Statistical average of student readout norms squared as a function of network width from theory and HMC sampling, for student-teacher tasks described in Section 4.2. (a): $\langle\|a_{l}\|^{2}\rangle\sigma_{w}^{2}/N$ as a function of network width $N$ for a fixed $\sigma_{w}$. The branch norms break the GP symmetry as it goes to the narrow width limit. (b)(c): Branch 0 and branch 1 readout norm squared respectively for a range of $\sigma_{w}$ regularization values. The student branch norms with different regularization strengths all converge to the same teacher readout norm values at narrow width.
@@
We perform theoretical calculations and HMC sampling for the student branch squared readout norms as we vary the student network width $N$ and the prior regularization strength $\sigma_{w}$. Figure 2(a) shows the statistical average of the student branch squared readout norms, ie. $\langle\|a_{l}\|^{2}\rangle\sigma_{w}^{2}/N$ as a function of the network width $N$, where the branch norms split as the network width gets smaller, which we call symmetry breaking. The symmetry breaking of branch norms from the GP limit to the narrow width limit accompanies the convergence to learning teacher’s norms at narrow width for different $\sigma_{w}$’s as shown in Figure 2(b)(c), supporting Eq. 21.
@@
We can also determine the generalization properties of the student readout lables at both the branch level $f_{l}$ (Appendix A.3) and overall level $f$ using Eq. 1819. We perform theoretical calculations and HMC sampling of generalization errors as a function of the network width $N$ and regularization strength $\sigma_{w}$, with results shown in Figure 3 and 4. At narrow width, we expect individual branch $f_{l}$ to learn the corresponding teacher’s branch output $f_{l}^{*}$ independently, causing the bias to increase with network width. This is observed for both branches, with a transition from the narrow-width regime to the GP regime. The regularization strength $\sigma_{w}$ controls the transition window, with larger $\sigma_{w}$’s leading to sharper transitions. This aligns with our analysis of the entropic and energetic contributions, where the larger $\sigma_{w}$ amplifies the distinction between the two terms. In contrast, the variance decreases with network width for small $\sigma_{w}$’s, resulting in a trade-off between the contributions of bias and variance to overall generalization performance, as shown clearly with the graph of network generalization vs. $N$ with $\sigma_{w}=0.5$ in Figure 4.
Figure 3: Student network squared bias and variance for individual branches as a function of network width $N$ and regularization strength $\sigma_{w}$. The mean and variance of branch $l$ readout $f_{l}^{\mu}$ for node $\mu$ is calculated in A.3 and the bias and variance for branch $l$ can be infered similarly as Eq. 19. Generalization values are normalized over the average true readout labels.
Figure 4: Student network generalization performance as a function of network width $N$ and regularization strength $\sigma_{w}$. Generalization is normalized over the average true readout labels.
@@ 5 BPB-GNN on Cora
We also perform experiments on the Cora benchmark dataset (McCallum et al. (2000)) by training the BPB-GNN with binary node regression, for a range of $L,N,\sigma_{w}$ values (Appendix B.2 for details). We observe a similar narrow-to-wide width transition for the bias term. As shown in Figure 5, the bias increases with network width, transitioning to the GP regime, and we observe the trend extending to a potential narrow width limit. [^2] In this case, the narrow width limit is hard to demonstrate as the transition window is below realistic minimum of network width. Additionally, it is demonstrated that using more branches that involve higher-order convolutions improves performance.
Figure 5: Cora generalization performance vs. network width $N$ and branch number $L$, for various regularization strength $\sigma_{w}$’s. The accuracy is computed by turning the mean predictor from HMC samples into a class label using its sign.
Convergence of branch readout norms at narrow width
An interesting aspect of the BPB-GNN network is that the branch readout norms converge at the narrow width for different hyperparameters $\sigma_{w}$ and $L$, reflecting the natural branch importance for the task.
Figure 6: Cora experiment: statistical average of squared readout norms $\langle\|a_{l}\|^{2}\rangle\sigma_{w}^{2}/N$ for each branch $l$ as a function of the network width $N$, and regularization strength $\sigma_{w}$. The BPB-GNN has $L=6$ branches.
As shown in Figure 6, the BPB-GNN with branches $L=6$ robustly learns the readout norms at narrow width independently of $\sigma_{w}$’ s, consistent with the student-teacher results. This suggests that we can recast the data as generated from a ground-truth teacher network even for real-world datasets. The last branch of the BPB-GNN network has a larger contribution, reflecting the presence of higher-order convolutions in the Cora dataset. From a kernel perspective, increasing branches better distinguish the nodes, as shown in Figure 7. This could explain the selective turn-off of intermediate branches and the increased contribution of the last branch.
Figure 7: Kernel $K_{l}=\frac{\sigma_{w}^{2}}{N_{0}}A^{l}X_{0}X_{0}^{T}A^{l}$ for each branch $l$ shown for the first 8 branches on Cora dataset, sorted by node labels. $A$ is the normalized adjacency matrix of the Cora graph, $X_{0}$ the node feature matrix and $N_{0}$ the node feature dimension. Initialization variance $\sigma^{2}_{w}=1$ and total node number $n=2708$.
Figure 8: Branch Importance vs. branch number $l$ on Cora. Legends represent different total number of branches $L$. The branch importance is defined as the statistical average of branch squared readout norms $\langle\|a_{l}\|^{2}\rangle\sigma_{w}^{2}/N$ at the narrow width limit; here we take the empirical branch norm values at $N=4$ and fixed $\sigma_{w}=1.4$.
Furthermore, the first two branches are learned most robustly at narrow width, as shown in Figure 8, where the branch norms converge for the first two branches even for BPB-GNNs with different $L$. This suggests that the branch importance, as reflected by the norms learned at narrow width, indicates the contribution of the bare data and the first convolution layer.
@@ 6 Discussion
The findings presented in this paper reveal that BPB-GNNs exhibit unique characteristics in the narrow width limit. Unlike the infinite-width limit, where neural networks behave as Gaussian Process (GP) with task-independent kernels, narrow-width BPB-GNNs undergo significant kernel renormalization. This renormalization leads to breaking of the symmetry between the branches, resulting in more robust and differentiated learning. Our experiments demonstrate that narrow-width BPB-GNNs can retain and, in some cases, improve generalization performance compared to their wider counterparts, particularly in bias-limited scenarios where the regularization effects dominate. Additionally, the observed independence of readout norms from architectural hyperparameters suggests that narrow-width BPB-GNNs can capture the intrinsic properties of the data more effectively. These insights suggest potential new strategies for optimizing neural network architectures in practical applications, challenging the traditional emphasis on increasing network width for better performance. Although our work is focused on GNN architectures, our findings can be transferred to the transformer family of architectures, as the transformer with frozen attention is analogous to a fully connected GNN (Vaswani et al. (2017); Veličković et al. (2017)), and a Bayesian theory with kernel renormalization can be developed similarly as in Tiberi et al. (2024).
@@ 7 Conclusion
In conclusion, this paper introduces and investigates the concept of narrow width limits in Bayesian Parallel Branching Graph Neural Networks. Contrary to the common belief that wider networks inherently generalize better, our results indicate that BPB-GNNs with significantly narrower widths can achieve better or competitive performance. This is attributed to effective symmetry breaking and kernel renormalization in the narrow-width limit, which leads to robust learning. Our theoretical analysis, supported by empirical evidence, establishes a new understanding of how network architecture influences learning outcomes. This work provides a novel perspective on neural network design and suggests further research into optimizing network structures for specific learning tasks, especially in scenarios where model simplicity and computational efficiency are crucial.
acknowledgement
@@ Appendix A Details on Theory of BPB-GNN
A.1 Summary of Notations
Hyperparameters and Dimensions
$P$ Number of training nodes
$n$ Total number of nodes for a graph
$N_{0}$ Input node feature dimension
$N$ BPB-GNN hidden layer width
$L$ Total number of branches
$\sigma_{w}$ $L_{2}$ prior regularization strength
$T$ Temperature
$\alpha_{0}=\frac{P}{LN_{0}}$ Network capacity
$\alpha=\frac{P}{N}$ Width ratio
Network Architecture and Input/Output
$\hat{A}\in\mathbb{R}^{n\times n}$ Adjacency matrix
$A\in\mathbb{R}^{n\times n}$ Normalized adjacency matrix by its degree matrix $D$
$X\in\mathbb{R}^{n\times N_{0}}$ Input node feature matrix
$G=(X,A)$ Graph
$W^{(l)}\in\mathbb{R}^{N_{0}\times N}$ Hidden layer weight for branch $l$
$a^{(l)},a_{l}\in\mathbb{R}^{N}$ Readout vector for branch $l$
$\Theta_{l}=(W_{l},a_{l})$ Collection of parameters for branch $l$
$h_{l}^{\mu}\in\mathbb{R}^{N}$ Activation vector for branch $l$ and node $\mu$
$f_{l}^{\mu}\in\mathbb{R}$ Readout prediction for branch $l$ and node $\mu$
$f^{\mu}\in\mathbb{R}$ Overall readout prediction for node $\mu$
$y^{\mu}\in\mathbb{R}$ Overall node label for node $\mu$
$H_{l}(W_{l})\in\mathbb{R}^{P\times N}$ Activation feature matrix for branch $l$
$Y\in\mathbb{R}^{P}$ Training node labels
$F\in\mathbb{R}^{P}$ Readout predictions
Statistical Theory
$E(\Theta;G,Y)$ Energy loss function
$Z$ Partition function
$I\in\mathbb{R}^{P\times P}$ Identity matrix
$H(W)$ Hamiltonian after integrating out readout $a_{l}$’s
$u\in\mathbb{R}^{L}$ Order parameters as saddle point solution
$u_{l}\in\mathbb{R}$ Order parameter for branch $l$
$H(u)$ Hamiltonian as a function of order parameters
$S(u)$ Entropy as a function of order parameters
$E(u)$ Energy as a function of order parameters
$r_{l}$ Mean squared readout
$\mathrm{Tr}_{l}$ Variance-related of readout
Kernels
$K(W)\in\mathbb{R}^{P\times P}$ Hidden layer weight dependent overall kernel
$K_{l}$ Branch $l$ kernel
$K\in\mathbb{R}^{P\times P}$ Overall kernel averaged over $W$’s, $K=\sum_{l}\frac{u_{l}K_{l}}{L}$
$k_{l}^{\nu}\in\mathbb{R}^{P\times 1}$ Branch $l$ kernel column for the $P$ training nodes against test node $\nu$
$k^{\nu}\in\mathbb{R}^{P\times 1}$ Overall kernel column for the $P$ training nodes against test node $\nu$
$|_{P}$ Kernel restricted to the $P$ training nodes against $P$ training nodes
$|_{(P,\nu)}$ Kernel restricted to $P$ training nodes against the test node $\nu$
$|_{(\nu,\nu)}$ Kernel restricted to the test node $\nu$ against test node $\nu$
Student-teacher Setup
$Y_{l}^{*}\in\mathbb{R}^{P}$ Teacher network readout prediction for branch $l$ for $P$ nodes
$Y^{*}\in\mathbb{R}^{P}$ Teacher network overall readout for $P$ nodes
$W_{l}^{*}\in\mathbb{R}^{N_{0}\times N}$ Teacher hidden layer weight for branch $l$
$A_{l}^{*}\in\mathbb{R}^{N}$ Teacher readout vector for branch $l$
$\beta_{l}^{2}$ Teacher readout variance
$\sigma_{t}^{2}$ Teacher hidden layer weight variance
A.2 Kernel renormalization
@@
$H(W)=\frac{1}{2\sigma_{w}^{2}}\sum_{l=0}^{L-1}\mathrm{Tr}W_{l}^{T}W_{l}+\frac{1}{2}Y^{T}(K(W)+TI)^{-1}Y+\frac{1}{2}\log\det(K(W)+TI),$ (27)
@@
$K(W)=\frac{1}{L}\sum_{l}\frac{\sigma_{w}^{2}}{N}(H_{l}(W_{l})H_{l}(W_{l})^{T})|_{P}$ (28)
@@
$Z$ $=\int\prod_{l=0}^{L}dh_{l}du_{l}dt\exp\Big{(}it^{T}Y-\sum_{l}\frac{N}{2}\log(1+h_{l})+\sum_{l}\frac{N}{2\sigma_{w}^{2}}u_{l}h_{l}-\frac{1}{2}t^{T}\Big{(}\sum_{l}\frac{1}{L}u_{l}K_{l}+TI\Big{)}t\Big{)}$
$=\int\prod_{l=0}^{L}dh_{l}du_{l}\exp\Big{(}-\sum_{l}\frac{N}{2}\log(1+h_{l})+\sum_{l}\frac{N}{2\sigma_{w}^{2}}u_{l}h_{l}\frac{1}{2}Y^{T}\Big{(}\sum_{l}\frac{1}{L}u_{l}K_{l}+TI\Big{)}^{-1}Y\Big{)}$ (29)
@@
$H_{eff}(u)=S(u)+E(u),$ (33)
where we call $S(u)$ the entropic term
$S(u)=-\sum_{l}\frac{N}{2}\log u_{l}+\sum_{l}\frac{N}{2\sigma_{w}^{2}}u_{l}$ (34)
and $E(u)$ the energetic term
$E(u)=\frac{1}{2}Y^{T}(\sum_{l}\frac{1}{L}u_{l}K_{l}+TI)^{-1}Y+\frac{1}{2}\log\det(\sum_{l}\frac{1}{L}u_{l}K_{l}+TI)$ (35)
@@
$N(1-\frac{u_{l}}{\sigma_{w}^{2}})=-Y^{T}(K+TI)^{-1}\frac{u_{l}K_{l}}{L}(K+TI)^{-1}Y+\mathrm{Tr}[K^{-1}\frac{u_{l}K_{l}}{L}],$ (38)
@@
$N(1-\frac{u_{l}}{\sigma_{w}^{2}})=-Y^{T}K^{-1}\frac{u_{l}K_{l}}{L}K^{-1}Y+\mathrm{Tr}[K^{-1}\frac{u_{l}K_{l}}{L}]$ (41)
A.3 Predictor statistics and generalization
@@
$\begin{split}Z(\eta_{1},\dots,\eta_{L})=\int D\Theta\exp{\left\{-\frac{\beta}{2}\sum_{\mu}(f^{\mu}(G;\Theta)-y^{\mu})^{2}\right.}\\ \left.+\sum_{l}i\eta_{l}\frac{1}{\sqrt{NL}}\sum_{i}a^{(l)}_{i}h_{i}^{(l),\nu}(G,W_{l})-\frac{T}{2\sigma_{w}^{2}}\Theta^{T}\Theta\right\}\end{split}$ (42)
@@
$\langle f_{l}^{\nu}(G)\rangle=\partial_{i\eta_{l}}\log Z\big{|}_{\vec{\eta}=0}$ (43)
$\langle\delta f_{l,\nu}^{2}(G)\rangle=\partial^{2}_{i\eta_{l}}\log Z\big{|}_{\vec{\eta}=0}$ (44)
@@
$Z(\eta_{1},\dots,\eta_{L})$ $=\int\Pi_{l}du_{l}\,\exp{\left\{\sum_{l}\left(\frac{N}{2}\log u_{l}-\frac{N}{2\sigma_{w}^{2}}u_{l}\right)\right.}$
@@
$\langle f_{l}^{\nu}\rangle=\frac{u_{l}k_{l,\nu}^{T}}{L}(K+TI)^{-1}Y$ (48)
@@
$\langle\delta f_{l,\nu}^{2}\rangle=\frac{u_{l}K_{l}^{\nu,\nu}}{L}-\frac{u_{l}k_{l,\nu}^{T}}{L}(K+TI)^{-1}\frac{u_{l}k_{l,\nu}}{L}$ (49)
The predictor statistics of the overall readout $f=\sum_{l}f_{l}$ is given by:
$\langle f^{\nu}(G)\rangle=\sum_{l}\frac{u_{l}k_{l,\nu}^{T}}{L}(K+TI)^{-1}Y=k_{\nu}^{T}(K+TI)^{-1}Y$ (50)
$\langle\delta f(G)_{\nu}^{2}\rangle=\sum_{l}u_{l}K_{l}^{\nu,\nu}-\sum_{l,l^{\prime}}u_{l}k_{l,\nu}^{T}(K+TI)^{-1}u_{l^{\prime}}k_{l^{\prime},\nu}=K_{\nu,\nu}-k_{\nu}^{T}(K+TI)^{-1}k_{\nu}$ (51)
A.4 Statistics of branch readout norms